OpenStudy (anonymous):

At Cheryl's Canine Resort, there are reservations for twice as many retrievers as collies, and for two more shepards than collies. If there are 21 cages in all, what is the greatest number of collies that can be boarded

OpenStudy (anonymous):

Let x be the number of collies. Let 2x be the number of retrievers. Let x+2 be the number of shepards. set x+2x+x+2=21 and solve

OpenStudy (anonymous):

you're going to get a fraction, but make sure to round down because you can't have part of a dog

OpenStudy (anonymous):

i need it to be an inequality so how would i do that??

OpenStudy (anonymous):

ok, so the total number of dogs (the left hand side of the equation) needs to be less than or equal to 21 (the right hand side of the equation) do you know the greater than or equal to symbol?

OpenStudy (anonymous):

or the less than or equal to symbol?

OpenStudy (anonymous):

Yea i do

OpenStudy (anonymous):

ok, so the number of dogs has to be less than or equal to 21. insert the proper symbol where I put the = sign and tell me what you think it is

OpenStudy (anonymous):

so is the answer x is greater than or equal to -9.5 or -19/2?

OpenStudy (anonymous):

\[x+2x+x+2 \le 21\]

OpenStudy (anonymous):

Ok, let's take a step back, I think you just made an arithmetic mistake

OpenStudy (anonymous):

is it x is less than 19/4?

OpenStudy (anonymous):

Yes perfect! Now what would that decimal be?

OpenStudy (anonymous):

*less than or equal to

OpenStudy (anonymous):

4.75?

OpenStudy (anonymous):

Ok, so we can have \[x \le 4.75\]

OpenStudy (anonymous):

but there cannot be .75 of a dog. So we have to round down

OpenStudy (anonymous):

so then its just 4?

OpenStudy (anonymous):

Correct. So 4 is our x value and x is our number of Collies. Therefore 4 is the maximum number of Collies that can be boarded. Does this make sense?

OpenStudy (anonymous):

yes it does thank you

OpenStudy (anonymous):

You're welcome :)